# Peter Reinhard Hansen§    and   Chen Tong

### v3

@@ Abstract
We introduce a novel method for obtaining a wide variety of moments of a random variable with a well-defined moment-generating function (MGF). We derive new expressions for fractional moments and fractional absolute moments, both central and non-central moments. The new moment expressions are relatively simple integrals that involve the MGF, but do not require its derivatives. We label the new method CMGF because it uses a complex extension of the MGF and can be used to obtain complex moments. We illustrate the new method with three applications where the MGF is available in closed-form, while the corresponding densities and the derivatives of the MGF are either unavailable or very difficult to obtain.
@@ 1 Introduction
Moments of random variables, including conditional moments, are central to much of economics, finance, and econometrics. In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF). The proposed method is general, computationally efficient, and can offer solutions to problems where standard methods are inadequate.
@@
It is well known that the $k$-th integer moment of a random variable $X$, with MGF, $M_{X}(s)=\mathbb{E}[e^{sX}]$, is given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ is the $k$-th derivative of $M_{X}(s)$. Non-integer moments and fractional absolute moments, $\mathbb{E}|X^{r}|$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF). [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022). For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), and Schürger, (2002). In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments. Unlike existing expressions, the new expressions do not involve derivatives (of the MGF or the CF). Thus, the new expressions are particularly useful when derivatives are prohibitively difficult to obtain. This is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions. The new method is labelled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$. Although complex moments are not commonly used in econometrics, this generalization can be included with minimal adaptation.  [^2] Complex moments are commonly used other fields, including statistical mechanics, number theory, and quantum physics. In part, because complex moments are related to Mellin transforms. The new expressions can also be used for establishing new identities, such as the expression for the reciprocal Gamma function, which we infer from the absolute moment of a Gaussian random variable.
@@
The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in most cases, they facilitate new ways to compute moments numerically, which are fast and accurate. We demonstrate this with the normal-inverse Gaussian distribution. We also show that the CMGF method offers fast and accurate moments in dynamic models, where we are not aware of good alternative methods. Specifically, we use the CMGF method to compute moments of cumulative returns in the Heston-Nandi GARCH (HNG) model, to compute moments of realized volatilities in the Heterogeneous Autoregressive Gamma (HARG) model, and to compute conditional moments in the Autoregressive Poisson Model.
@@
The remainder of this paper is organized as follows. We present the new moment expressions in Section 2. Theorem 1 has the expression for absolute fractional moments for general random variables, Theorem 2 has the expression for fractional moments of non-negative random variables, and Theorem 3 presents a new expression for integer moments. As an example, we apply CMGF to the normal-inverse Gaussian distribution. Section 3 presents three applications of the new expressions in dynamic models, and Section 4 concludes.
@@ 2 A New Method for Computing Moments (CMGF)
We will use the following notation: $\mathbb{N}$ denotes the positive integers and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. We use $\mathcal{N}$ to denote a neighborhood about zero, for which $(-\delta,\delta)\subset\mathcal{N}$ for some $\delta>0$, and the positive part of $\mathcal{N}$ is denoted $\mathcal{N}_{+}\equiv\{s\in\mathcal{N}:s>0\}$. We use $z=s+it$ and $\zeta=\sigma+i\tau$ to represent complex variables, where $s=\operatorname{Re}(z)$ and $t=\operatorname{Im}(z)$, and similarly for $\zeta$.
To avoid possible confusion, we use $\mathbb{E}[X^{k}]$ to denotes integer moments, $k\in\mathbb{N}$, whereas general moments (including fractional and complex moments) are represented with $\mathbb{E}[X^{r}]$, $r\in\mathbb{C}$.
Our results are applicable to any random variable with a well-defined MGF, as characterized in the following assumption.
Assumption 1.
$X$ is a real random variable with
$M_{X}(s)=\mathbb{E}[e^{sX}]<\infty,\qquad\text{for all}\quad s\in\mathcal{N},$
where $\mathcal{N}$ is a neighborhood about zero.
The MGF is typically used as a real function with a real argument, such that both $s$ and $M_{X}(s)$ take values in $\mathbb{R}.$ In this paper, we establish our results by means of complex arguments, $z\in\mathbb{C},$ such that $M_{X}(z)$ also takes values in $\mathbb{C}$.
@@
nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases. The complex valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform.  [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent. Moreover, $M_{X}(z)$ is obviously also applicable to discrete distributions and distributions that are mixtures of continuous and discrete distributions.
The results are based on the following identities.
@@
Let $x\in\mathbb{R}$ and $z=s+it\in\mathbb{C}$, where $s>0$ is an arbitrary positive constant.
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1,$
for any $x\neq 0$, and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$. For $x>0$ we have
$x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1.$
@@
$x^{k}$ $=$ $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,\qquad k\in\mathbb{N}_{0},$
for any $x\neq 0$, and this identity also holds for $x=0$ if $k\in\mathbb{N}$.
The case, $x=0$, arises as a special case in Lemma 1, requiring results to include a qualification for distributions with positive mass at zero.
Our first result is the following expression for fractional absolute (possibly complex) moments.
@@
Suppose $X$ satisfies Assumption 1, $\xi\in\mathbb{R}$, and $z=s+it$ with $s\in\mathcal{N}_{+}$. Then
$\mathbb{E}|X-\xi|^{r}$ $=$ $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\mathrm{d}t,$ (1)
for $\operatorname{Re}(r)\geq 0$, and if $\Pr(X=\xi)=0$, then the identity holds for all $\operatorname{Re}(r)>-1$.
If $r$ is real, then (1) is equivalent to
$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\right]\mathrm{d}t.$ (2)
The result obviously covers regular and central absolute moments by setting $\xi=0$ for regular moments and $\xi=\mathbb{E}X$ for central moments. Note that the expression remains invariant to the specific value of $s$, so long as $s\in\mathcal{N}_{+}$. In our experience, the choice of $s$ has no practical impact on moment evaluation. Although analytical integration may often be impractical, (1) offers a novel method to evaluate moments numerically. For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (2) rather than (1).
Interestingly, Theorem 1 can be used to obtain new identities, by equating (1) with an existing expression for absolute moments. For instance, the absolute moment for standard normal random variable, $X\sim N(0,1)$, is $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for $r>-1$, and equating this with (1) yields the following expression for the reciprocal gamma function
@@
for $z=s+it$, where $s>0$ is an arbitrary positive constant. While we can only infer that (3) holds for real $r>-1$, the identity actually holds for any $r\in\mathbb{C}$, see Hansen and Tong, (2024).
For non-negative random variables we have the following simpler expression for fractional moments.
Theorem 2 (Moments of non-negative random variables).
Let $X$ satisfies Assumption 1 and $\Pr(X\geq\xi)=1$ for some $\xi\in\mathbb{R}$. For any $r\in\mathbb{C}$ with $\operatorname{Re}(r)\geq 0$, we have,
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,$ (4)
where $z=s+it$ with $s\in\mathcal{N}_{+}$. If $X$ is strictly greater than $\xi$, i.e. $\Pr(X>\xi)=1$, then the identity holds for $\operatorname{Re}(r)>-1$.
If $r$ is real, then (4) is equivalent to
@@
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t.$ (5)
@@
For an exponentially distributed random variable, with parameter $\lambda>0$, $X\sim\operatorname{Exp}(\lambda)$, we have $M_{X}(z)=\frac{\lambda}{\lambda-z}$, with $\mathcal{N}=(-\infty,\lambda)$. So, the integral in (5) is $\int_{0}^{\infty}{\rm Re}\left[\frac{\lambda}{\lambda-z}\frac{1}{z^{r+1}}\right]\mathrm{d}t$. For an integer moment, $k\in\mathbb{N}$, this integral equals
@@
because $\int_{0}^{\infty}{\rm Re}\left[z^{-j}\right]\mathrm{d}t=0$ for $j>1$. Substituting $z=s+it$ in the last expression gives us
@@
Thus that $\mathbb{E}[X^{k}]=\lambda^{-k}k!$ as expected. Deriving non-integer moments, $r\notin\mathbb{N}$, can be derived using contour integrals and the Cauchy Integral Theorem. These are the techniques used extensively to prove the main results in this paper.

### v4

@@ Abstract
We introduce a novel method for obtaining a wide variety of moments of any random variable with a well-defined moment-generating function (MGF). We derive new expressions for fractional moments and fractional absolute moments, both central and non-central moments. The expressions are relatively simple integrals that involve the MGF, but do not require its derivatives. We label the new method CMGF because it uses a complex extension of the MGF and can be used to obtain complex moments. We illustrate the new method with three applications where the MGF is available in closed-form, while the corresponding densities and the derivatives of the MGF are either unavailable or very difficult to obtain.
@@ 1 Introduction
Moments of random variables, including conditional moments, are central to many areas of hard and social sciences. In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF). The proposed method is general, computationally efficient, and can offer solutions to problems where existing methods are inadequate.
@@
It is well known that the $k$-th integer moment of a random variable $X$, with MGF, $M_{X}(s)=\mathbb{E}[e^{sX}]$, is given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ is the $k$-th derivative of $M_{X}(s)$. Non-integer moments and fractional absolute moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF). [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022). For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), Schürger, (2002), and Meng, (2005). In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments. The new expressions stand out by being applicable to all random variables with a MGF and can be used to compute moments of $Y=g(X)$, whenever $g(x)$ can be expressed as a linear combination of a finite number of fractional, integer, absolute, positive-part, and possibly centralized powers of $x$. The new expressions are particularly useful when derivatives are prohibitively difficult to obtain, as is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions. The new method is also useful in situations where the density is unavailable, as can be the case for compound distributions, exponential tilting, and some Lévy processes.
The new method is labeled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$ with $\operatorname{Re}(r)>-1$. Although complex moments are not commonly used in econometrics, this generalization can be included with minimal adaptation. [^2] Complex moments are commonly used other fields, including quantum physics, number theory, and statistical mechanics.
@@
The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in most cases, they facilitate new ways to compute moments numerically, which are fast and accurate. We demonstrate this with the normal-inverse Gaussian (NIG) distribution. Existing expressions for absolute fractional moments of the NIG distribution are limited to $\mu$-centered moments, and the expressions are infinite series involving modified Bessel functions of the second kind. The CMGF integral expressions are simpler and do not involve special functions. We also show that the CMGF method offers fast and accurate moments in dynamic models, where we are not aware of good alternative methods. Specifically, we use the CMGF method to compute moments of cumulative returns in the Heston-Nandi GARCH (HNG) model, to compute moments of realized volatilities in the Heterogeneous Autoregressive Gamma (HARG) model, and to compute conditional moments in the Autoregressive Poisson Model.
A key step in our proof is the identity, $\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t=0$, that holds for $s>0$, $x>0$ and $\operatorname{Re}(r)>-1$. This identity simplifies expressions involving an inverse Laplace transform, such that the moment-generating function emerges. The identity, see Lemma 1, appears to be new, which can explain that these simple and general integral expressions were not discovered earlier. We are not the first to derive integral expressions that involve the MGF. Meng, (2005) derived integral expressions of $\mathbb{E}[X^{a}/Y^{b}]$, involving their joint MGF $M_{X,Y}(s_{1},s_{2})$ and its derivatives. The possibility of combining characteristic and moment-generating functions, $\mathbb{E}[e^{sY+itX}]$, is mentioned in the epilogue of Meng, (2005). For $X=Y$ this is an extension of the MGF (or the CF) to the complex plane, $M_{X}(s+it)$, which is at the core of our expression.
An auxiliary result that arises from applying the integral expression to standard normal distribution, is a holomorphic integral representation of Gamma function and its reciprocal, which do not rely on analytic continuation, see Hansen and Tong, (2025).
@@
The remainder of this paper is organized as follows. We present the new moment expressions in Section 2. Theorem 1 has the expression for absolute fractional moments for general random variables, Theorem 2 presents positive-part moments, and Theorem 3 presents a new expression for integer moments. Corollary 1 has expressions for fractional moments of nonnegative (positive) random variables and Corollary 2 has expression for (odd integers) negative tail moments. We apply CMGF to the normal-inverse Gaussian distribution. Section 3 presents three applications of the new expressions in dynamic models, and Section 4 concludes.
@@ 2 A New Method for Computing Moments (CMGF)
We use the following notation: $\mathbb{N}$ denotes the positive integers and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. We use $z=s+it\in\mathbb{C}$, where $s=\operatorname{Re}(z)>0$. We use $\mathbb{E}[X^{k}]$ to denote integer moments, $k\in\mathbb{N}_{0}$, whereas $\mathbb{E}[X^{r}]$, denotes general moments including fractional and complex moments, $r\in\mathbb{C}$. The MGF is typically introduced as a real function with a real argument, such that both $s$ and $M_{X}(s)$ take values in $\mathbb{R}.$ In this paper, we establish our results by means of complex arguments, $z\in\mathbb{C},$ such that $M_{X}(z)$ also takes values in $\mathbb{C}$.
Interpretation and scope: For complex powers $z^{-(r+1)}$, we fix the principal branch of the complex logarithm (branch cut on $(-\infty,0]$) and interpret integrals as oscillatory (i.e., symmetric Cauchy principal values) whenever $1<\operatorname{Re}(r)\leq 0$. Assumptions on one-sided exponential moments of $X$ (i.e., $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$) are stated with each theorem. When $\operatorname{Re}(r)\in(-1,0)$, we additionally assume $\mathbb{E},\lvert X-\xi\rvert^{r}<\infty$ to justify exchanging expectation and integration.
@@
nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases. The complex-valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform. [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent. Moreover, $M_{X}(z)$ is applicable to discrete distributions and mixtures of continuous and discrete distributions.
The following result is key to several simplifications.
@@
$\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t=0,\qquad\text{for}\quad x>0,\ s>0,\ \operatorname{Re}(r)>-1.$ (1)
For $x=0$ the identity holds for $\operatorname{Re}(r)>0$. [^4] Obviously, for $x<0$ the identity is $\int_{-\infty}^{+\infty}\frac{e^{itx}}{(s+it)^{r+1}}\mathrm{d}t=0$ for $s>0$ and $\operatorname{Re}(r)>-1$.
The identity in (1) is obtained by a contour argument that is valid for $\operatorname{Re}(r)>-1$. Note that $f\in L^{1}(\mathbb{R})$ if and only if $\operatorname{Re}(r)>0$, because
$\left|f(t)\right|=\left|\frac{e^{-itx}}{(s+it)^{r+1}}\right|=\frac{1}{|s+it|^{\operatorname{Re}(r)+1}}=\frac{1}{(s^{2}+t^{2})^{(\operatorname{Re}(r)+1)/2}}.$
For $1<\operatorname{Re}(r)\leq 0$ the integral is not absolutely convergent; in this range we interpret the formula (1) as the oscillatory improper integral,
$\lim_{T\to\infty}\int_{-T}^{T}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t,$
which exists and equals 0 for $x>0$ by closing the contour in the lower half-plane ($e^{-itx}$ decays there) and observing that $s+it\neq 0$ for $t\in\mathbb{R}$ when $s>0$.
For the special case $x=0$ and $r=0$, it is easy to establish the well-known identity
$\int_{-\infty}^{+\infty}\frac{1}{s+it}\mathrm{d}t=\pi\operatorname{sgn}(s),\qquad\text{for all }s\in\mathbb{R},$ (2)
where $\operatorname{sgn}(s)\equiv\frac{s}{|s|}1_{\{s\neq 0\}}\in\{-1,0,1\}$. In this setting where $s>0$, we simply have $\int_{-\infty}^{+\infty}\frac{1}{s+it}\mathrm{d}t=\pi$, which we will use in the first formula of following Lemma.
Lemma 2.
Let $x\in\mathbb{R}$ and $z=s+it\in\mathbb{C}$, where $s>0$ is an arbitrary constant. Then
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for}\quad\operatorname{Re}(r)>-1,\quad x\neq 0,$
and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$ or r=0r=0.555When $x^{r}$ is to be evaluated for $x=r=0$, we follow the standard convention $0^{0}=1$.
For $x>0$ we have
$x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for}\quad\operatorname{Re}(r)>-1,\quad x>0,$
and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$.
@@
$x^{k}$ $=$ $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,\qquad k\in\mathbb{N}_{0},\quad x\in\mathbb{R}.$
Here $x=0$, arises as a special case in the first two formulas of Lemma 2, and this requiring results to include a qualification for distributions with positive mass at zero.
Several results will now follows by taking expectation to both sides of the identities in Lemma 2, and justify interchanging expectation and integral. For $\operatorname{Re}(r)>0$ this is straight forward by means of Fubini/Tonelli theorems. However, as detailed above, the integrand is not absolute integrable for $\operatorname{Re}(r)\leq 0$, and interchanging expectation and integral requires a different line of arguments, and the improper integrals in the form $\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$ is to be interpreted as $\lim_{T\rightarrow\infty}\int_{-T}^{+T}[\cdot]\mathrm{d}t$ for $\operatorname{Re}(r)\in(-1,0]$.
Our first main result concerns fractional absolute moments.
@@
Suppose that $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$. For $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. If $\Pr(X=\xi)=0$ then
$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\right]\mathrm{d}t,\quad\text{for }r>-1,$ (3)
and for complex moments, $r\in\mathbb{C}$, we have
$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\mathrm{d}t,\quad\operatorname{Re}(r)>-1.$ (4)
Moreover, if $\Pr(X=\xi)>0$, the identities hold for $r\geq 0$ and $\operatorname{Re}(r)>0$ respectively.
The finite moment requirement, $\mathbb{E}|X-\xi|^{r}<\infty$, is implied by $\mathbb{E}[e^{\pm sX}]<\infty$ for $\operatorname{Re}(r)>0$, but is not guaranteed for $\operatorname{Re}(r)<0$. The results in Theorem 1 include regular absolute moments by setting $\xi=0$ and central moments by setting $\xi=\mathbb{E}X$. Although analytical integration may often be impractical, (4) offers a novel method to evaluate moments numerically. For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (3) rather than (4).
Remark.
$\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$ is equivalent to assuming the MGF is well defined in a neighborhood of zero. In next Theorem we only assume $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$, which does not guarantee the MGF is well-defined in a neighborhood about zero.
Interestingly, Theorem 1 can be used to obtain new identities, by equating (4) with an existing expression for absolute moments. For instance, the absolute moment of a standard normal random variable, $X\sim N(0,1)$, is $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for r>1r>-1, and equating this with (4) yields the following expression for the reciprocal gamma function
@@
for $z=s+it$, where $s>0$ is an arbitrary positive constant. From Theorem 1 we can only infer that (5) holds for real r>1r>-1, but the identity actually holds for any $r\in\mathbb{C}$, such that we obtain a new holomorphic integral representation of Gamma function and its reciprocal that does not rely on analytic continuation, see Hansen and Tong, (2025).
Below we will derive an expression for positive-part moments, where we use the convention $x_{+}^{r}=x^{r}1_{\{x>0\}}$. This enables us to accommodate negative powers, because $0_{+}^{r}=0$ for all rr. [^6] Note that with this convention we have ($r=0$) we have $0=0_{+}^{0}$, unlike $0^{0}=1$). Importantly, for $\operatorname{Re}(r)\in(-1,0]$, the integral $\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t$, is to be interpreted as the symmetric Cauchy principal value, $\lim_{T\rightarrow\infty}\int_{-T}^{+T}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t$.
Theorem 2 (Positive-part moments).
Suppose that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$. For $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. Then
@@
$\mathbb{E}[(X-\xi)_{+}^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t,\quad\text{for }r>0,$ (6)
and the the identity also holds for $r\in(-1,0]$ if $\Pr(X=\xi)=0$.
For complex moments, $r\in\mathbb{C}$, we have
$\mathbb{E}[(X-\xi)_{+}^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,\quad\text{for }\operatorname{Re}(r)>-1.$ (7)
Note that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ ensures that $\mathbb{E}(X-\xi)_{+}^{r}<\infty$ for ${\rm Re}\left(r\right)>0$ but not for $\operatorname{Re}(r)\in\left(-1,0\right)$. Therefore, for $\operatorname{Re}(r)\in\left(-1,0\right)$ where we rely on the dominated convergence theorem, the condition $\mathbb{E}|X-\xi|^{r}<\infty$ is required for interchanging expectation and integral.
Remark 1.
Here we assume $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$, which does not guarantee the existence of the MGF in an open neighborhood of zero. So, $\mathbb{E}|X-\xi|^{r}<\infty$ is not guaranteed by $\mathbb{E}[e^{sX}]<\infty$, as illustrated with the example below. Note that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ does ensure $\mathbb{E}(X-\xi)_{+}^{r}<\infty$ for $\operatorname{Re}(r)>0$.
Example 1 (One-sided MGF and divergent moments.).
Let $Y\sim\mathrm{Exp}(1)$, $Z\sim\mathrm{Pareto}(\alpha)$ with $\mathbb{P}(Z>z)=z^{-\alpha}$ for $z\geq 1$, and choose $0<\alpha<1$. Define
$X=\begin{cases}+Y,&\text{with prob. }1/2,\\ -Z,&\text{with prob. }1/2,\end{cases}$
with $Y$, $Z$ independent. Then, for $0<s<1$, $\mathbb{E}[e^{sX}]=\tfrac{1}{2}\mathbb{E}[e^{sY}]+\tfrac{1}{2}\mathbb{E}[e^{-sZ}]=\tfrac{1}{2}(1-s)^{-1}+\tfrac{1}{2}\mathbb{E}[e^{-sZ}]<\infty$. However, for $\tilde{s}<0$ we have $\mathbb{E}[e^{\tilde{s}X}]\geq\tfrac{1}{2}\mathbb{E}[e^{|\tilde{s}|Z}]=\tfrac{1}{2}\int_{1}^{\infty}e^{|\tilde{s}|z}\alpha z^{-\alpha-1}\mathrm{d}z=\infty$. Hence the MGF is not finite on any open neighborhood of 0. For $r\in(-1,\alpha)$ we have $\mathbb{E}|X|^{r}=\tfrac{1}{2}\Gamma(r+1)+\tfrac{1}{2}\frac{\alpha}{\alpha-r}$, while for $r\geq\alpha$, $\mathbb{E}[|X|^{r}]\geq\tfrac{1}{2}\mathbb{E}[Z^{r}]=\infty$. We can nevertheless apply Theorem 2 for r>1r>-1 because $\mathbb{E}(X)_{+}^{r}=\tfrac{1}{2}\Gamma(r+1)$.
The expression in Theorem 2 is identical to one in Pinelis, (2011, theorem 1), albeit we extend the range of moments to include some negative moments, $r\in(-1,0)$, and complex moments with $\operatorname{Re}(r)>-1$. The special case $r=1$ has received considerable attention because $\mathbb{E}[(X-\xi)_{+}]$ plays a central role in option pricing, with $\xi$ representing the strike price. For this special case, $r=1$, (7) coincides with Kim et al., (2010, lemma 7.1) and (6) coincides with Huang and Oosterlee, (2011, eq. 3.1). [^7] They derived the result under slightly stronger assumptions. In Kim et al., (2010) it is assumed that the distribution is infinitely divisible and continuous, while Huang and Oosterlee, (2011, eq. 3.1) assume continuity of the distribution. Neither of these assumptions is required here.
Note that the expression remains invariant to the value of $s>0$, and numerical integration is, in our experience, insensitive to the choice of $s$, so long as $\mathbb{E}[e^{sX}]<\infty$. For the positive-part moments, we do not require $\mathbb{E}[e^{uX}]<\infty$ for any $u<0$.
For non-negative random variables the following is a simple implication of Theorem 2 and Lemma 2.
Corollary 1 (Moments of nonnegative random variables).
Suppose that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ and for $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. (Strictly positive). If $\Pr(X>\xi)=1$, then
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t\quad\text{for }r>-1.$ (8)
(Nonnegative). If $\Pr(X\geq\xi)=1$, then the identity holds for $r>0$, and for complex moments, $r\in\mathbb{C}$, we have the identity
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,\quad\text{for }\operatorname{Re}(r)>-1.$ (9)
The condition $\mathbb{E}[(X-\xi)^{r}]<\infty$ is redundant for $\operatorname{Re}(r)>0$, because it is implied by $\mathbb{E}[e^{sX}]<\infty$, but for negative moments this additional conditions is needed.
@@
For an exponentially distributed random variable, with parameter $\lambda>0$, $X\sim\operatorname{Exp}(\lambda)$, we have $M_{X}(z)=\frac{\lambda}{\lambda-z}$, with region of convergence $\operatorname{Re}(z)<\lambda$. So, the integral in (8) is $\int_{0}^{\infty}{\rm Re}\left[\frac{\lambda}{\lambda-z}\frac{1}{z^{r+1}}\right]\mathrm{d}t$. For an integer moment, $k\in\mathbb{N}$, this integral equals
@@
as $\int_{0}^{\infty}{\rm Re}\left[z^{-j}\right]\mathrm{d}t=0$ for $j>1$. Substituting $z=s+it$ in the last expression gives us
@@
Thus that $\mathbb{E}[X^{k}]=\lambda^{-k}k!$ as expected. Non-integer moments, $r\notin\mathbb{N}$, can be derived using contour integrals and the Cauchy Integral Theorem.